### v2

# Time-dependent conserved operators for autonomous systems and quantization of resistance

Abstract
Two systems for a charged particle are studied, the first one when it is under the effect of a constant electric field, and the second one when it is under the effect of a constant electromagnetic field. For both systems, it is possible to find time-dependent conserved operators that can be used to derive time-dependent wave functions to the complete Schrödinger equation, such that the time variable is not separable from the space coordinates. At the same time, these conserved operators are used to build up a unitary operators, which define the symmetries of the systems. Then, it is shown that the invariance of the wave function under the action of these unitary operators leads to the quantization of resistance as integer multiples of the Klitzing’s constant.
Time-dependent operators, Resistance quantization, Klitzing’s constant
I Introduction
Conserved operators in dynamical systems have proven to be valuable tools in quantum mechanics for identifying system bases. For example, in the central force problem, angular momentum is conserved and used to determine the eigenfunctions associated with azimuthal and polar angles [1]. Another illustrative case involves a constant magnetic field, where, using Landau’s gauge, linear momentum is conserved and used to find the system’s basis. This is known as Landau’s ansatz, and the resulting basis turns out to be the product of plane waves and the harmonic oscillator wave function [2]. Recently, for a constant electromagnetic field, it has been shown that, in addition to Landau’s ansatz, a time-dependent conserved operator is necessary for a complete description of the system [3]. In the latter work mentioned, it was demonstrated that the invariance of the wave function under the action of a unitary transformation leads to the quantization of resistivity for a single particle, yielding a result similar to the experimentally measured expression by Störmer [4, 5, 6]. Furthermore, this invariance condition leads to two quantized quantities: the magnetic flux quantization (also derived using Landau’s ansatz [2]) and a second quantization condition related to the electric field, given by Eq. (34). The latter condition is the principal point of this work.
Two autonomous systems exhibiting time-dependent conserved operators are presented here. The first system involves a particle under the influence of a constant electric field, while the second describes a particle subjected to a constant electromagnetic field. The system with a linear potential is of special interest when studying the shift in energy levels (Stark effect) [7, 8, 9, 10] and time-dependent solutions have been found using the Baker-Campbell-Hausdorff generalized formula for arbitrary initial conditions [11]. However, the approach presented in this paper differs from the previous one by being based on the conserved properties of the system, also known as the Lewis and Riesenfeld invariant method. Although it was initially developed to address non-autonomous systems [12, 13], here it is shown that it can also be applied to autonomous systems. On the other hand, the second system when a constant magnetic field is added such that is parallel to the electric field had been used to study the photodetachment from negative ions and the spatial distribution of monoenergetic electrons [14, 15]. However, aside from these applications, the main result of this work is to demonstrate that these time-dependent wave functions require the quantization of resistance, which is proportional to an integer multiple of Klitzing’s constant [16, 17, 18], when they are invariant under the action of a unitary transform defined by the time-dependent conserved operator.
II About time-dependent conserved operators
The Hamiltonian we are interested in is written in CGS units as
${\hat{H}}=\frac{1}{2m}\left({\bf\hat{P}}-\frac{q}{c}{\bf A}\right)^{2}+qU,$ (1)
where $m$ is the mass of the particle, $q$ is the charge of the particle, $c$ is the speed of light, ${\bf\hat{P}}=({\hat{p}}_{x},{\hat{p}}_{y},{\hat{p}}_{z})=-i\hbar\nabla$ is the momentum operator, ${\bf A}=(A_{x},A_{y},A_{z})$ is the magnetic vector potential such that the magnetic field is given by its curl, that is ${\bf B}=\nabla\times{\bf A}$ and $U=U(x,y,z)$ is the electric potential such that the electric field is given by the negative of its gradient, ${\bf E}=-\nabla U$. Given our interest in identifying time-dependent conserved operators, it is important to demonstrate that they can be used to find solutions to the Schrödinger equation. The total variation of an operator, ${\hat{f}}$, is given by the Heisenberg equation
$\frac{d{\hat{f}}}{dt}=\frac{1}{i\hbar}[{\hat{f}},{\hat{H}}]+\frac{\partial{\hat{f}}}{\partial t}.$ (2)
When the above expression is equals to zero, one can say that the operator is conserved, that is
$\frac{1}{i\hbar}[{\hat{f}},{\hat{H}}]+\frac{\partial{\hat{f}}}{\partial t}=0.$ (3)
Here, ${\hat{f}}$ is a time-dependent operator which can be used to write down an eigenvalue equation
${\hat{f}}\psi=\lambda\psi,$ (4)
where $\lambda$ is a scalar. Then, applying Eq.(3) to $\psi$ we can write
${\hat{f}}({\hat{H}}\psi)-{\hat{H}}({\hat{f}}\psi)+i\hbar\frac{\partial{\hat{f}}}{\partial t}\psi=0,$ (5)
using Eq.(4) and $\partial_{t}({\hat{f}})\psi=\partial_{t}({\hat{f}}\psi)-{\hat{f}}\partial_{t}\psi$ the above expression can be rewritten as follows
$({\hat{f}}-\lambda)\left({\hat{H}}\psi-i\hbar\frac{\partial\psi}{\partial t}\right)=0.$ (6)
The above equality can be satisfied for the following reasons: 1) because $\hat{H}\psi-i\hbar\partial_{t}\psi$ is an element of the kernel of the operator $\hat{f}-\lambda$ defined as the set of all square-integrable functions in $\mathbb{R}^{3}$ that satisfies the previous expression,
$\text{Ker}(\hat{f}-\lambda)=\left\{\Phi\in L^{2}(\mathbb{R}^{3})|(\hat{f}-\lambda)\Phi=0\right\}.$ (7)
2) Because $\psi$ satisfies the Schrödinger equation
${\hat{H}}\psi=i\hbar\frac{\partial\psi}{\partial t}.$ (8)
III One-dimensional constant electric field
The simplest example where this approach can be used is when we work with a constant force in one dimension. In this case, the force is produced by the electric field, and the electric potential is given by $U(x)=-{\cal E}x$ where ${\cal E}$ is a constant representing the electric field intensity. The time-dependent Schrödinger equation is
$i\hbar\frac{\partial\psi}{\partial t}=\frac{1}{2m}{\hat{p}_{x}}^{2}\psi-q{\cal E}x\psi,$ (9)
for this case, it is known that the solutions for the stationary equation are the Airy function [19]. However, using Eq.(2) one can realize that this system has the following conserved operators
${\hat{f}}={\hat{p}}_{x}-q{\cal E}t,$ (10)
${\hat{E}}=i\hbar\frac{\partial}{\partial t}.$ (11)
Therefore one can write down the eigenvalue equation for the operator (10),
${\hat{p}}_{x}\psi-q{\cal E}t\psi=-q{\cal E}\delta t\psi,$ (12)
where, for simplicity, the eigenvalues were written as $-q{\cal E}\delta t$ such that $\delta t$ is a constant. Solving the above equation, it can be found that
$\psi(x,t)={\cal C}(t)\exp\left(i\frac{q{\cal E}}{\hbar}(t-\delta t)x\right),$ (13)
where ${\cal C}(t)$ is a function that depends only on time. To determine this function, we substitute the above wavefunction into Eq.(9) to obtain the following equation
$i\hbar\frac{\partial{\cal C}}{\partial t}=\frac{q^{2}{\cal E}^{2}}{2m}(t-\delta t)^{2}{\cal C},$ (14)
having the following solution
${\cal C}(t)=\exp\left(-i\frac{q^{2}{\cal E}^{2}}{6m\hbar}(t-\delta t)^{3}\right),$ (15)
and the solution can be written as
$\psi(x,t)=\frac{1}{L}\exp\left(-i\frac{q^{2}{\cal E}^{2}}{6m\hbar}(t-\delta t)^{3}+i\frac{q{\cal E}}{\hbar}(t-\delta t)x\right),$ (16)
where the normalization was calculated by integrating in a length of $x\in[-L/2,L/2]$. At this point, one can search for the simplest form of the solution, in a sense that the above wave function is just a unitary transform of the new function. The conserved operators in Eq.(10) and Eq.(11) can be used to define a couple of unitary operators
${\hat{U}}_{x}=\exp\left(-i\frac{\delta x}{\hbar}({\hat{p}}_{x}-q{\cal E}t)\right),$ (17)
${\hat{U}}_{t}=\exp\left(i\frac{\delta t}{\hbar}{\hat{E}}\right),$ (18)
where $\delta x$ is a constant. These operators define the symmetries of the system such that the Schrödinger equation remains invariant under a unitary transformation, i.e.,
${\hat{H}}-{\hat{E}}={\hat{U}}_{x}^{\dagger}({\hat{H}}-{\hat{E}}){\hat{U}}_{x}={\hat{U}}_{t}^{\dagger}({\hat{H}}-{\hat{E}}){\hat{U}}_{t}.$ (19)
Note that the wave function in Eq.(16) can be written as
$\psi(x,t)={\hat{U}}_{t}\frac{1}{L}\exp\left(-i\frac{q^{2}{\cal E}^{2}}{6m\hbar}t^{3}+i\frac{q{\cal E}}{\hbar}tx\right),$ (20)
which satisfies
${\hat{H}}\psi(x,t)-{\hat{E}}\psi(x,t)=0.$ (21)
multiplying the above equality by ${\hat{U}}_{t}^{\dagger}$, using Eq.(19) and the fact that ${\hat{U}}_{t}$ is unitary, we have
${\hat{H}}\varphi(x,t)-{\hat{E}}\varphi(x,t)=0,$ (22)
where
$\varphi(x,t)={\hat{U}}_{t}^{\dagger}\psi=\frac{1}{L}\exp\left(-i\frac{q^{2}{\cal E}^{2}}{6m\hbar}t^{3}+i\frac{q{\cal E}}{\hbar}tx\right).$ (23)
And this is the simplest form of the searched wave function. However, there is still another important characteristic of this solution to be analyzed, and is its degeneracy. Since the operator in Eq.(11) is conserved, $[{\hat{E}},{\hat{H}}]=0$, the following equality is satisfied
${\hat{E}}({\hat{H}}\varphi)={\hat{H}}({\hat{E}}\varphi),$ (24)
using the fact that $\varphi$ satisfies Eq.(9), the previous equality can be written as
$i\hbar\frac{\partial({\hat{E}}\varphi)}{\partial t}={\hat{H}}({\hat{E}}\varphi),$ (25)
it proves that
${\hat{E}}\varphi(x,t)=\left(\frac{q^{2}{\cal E}^{2}}{2m}t^{2}-q{\cal E}x\right)\varphi$ (26)
is another solution of Eq.(9). This characteristic can be generalized to any $j\in\mathbb{N}$ applications of the operator ${\hat{E}}$, since $[{\hat{E}}^{j},{\hat{H}}]=0$ one can prove that
$i\hbar\frac{\partial({\hat{E}}^{j}\varphi)}{\partial t}={\hat{H}}({\hat{E}}^{j}\varphi).$ (27)
Hence, we have a countable set of solutions. Defining each one of these functions as $\varphi^{j}(x,t)={\hat{E}}^{j}\varphi$ such that $\varphi^{0}=\varphi$, the following commutator can be calculated
$[{\hat{f}},{\hat{E}}^{j+1}]=i\hbar q{\cal E}(j+1){\hat{E}}^{j},$ (28)
and applying it to $\varphi$ we have that each one of these functions satisfies the following eigenvalue equation
${\hat{f}}{\hat{E}}\varphi^{j}=i\hbar q{\cal E}(j+1)\varphi^{j}.$ (29)
It is important to note that the eigenvalues for the above equation are imaginary. Despite the fact that the operators in Eq.(10) and Eq.(11) are Hermitian on their own, the product of two Hermitian operators is not necessarily Hermitian. Hence, ${\hat{f}}{\hat{E}}$ can have imaginary eigenvalues.
The general solution of this system can be written, as a superposition of the wave function Eq.(23) and its degeneration, that is
$\Psi(x,t)=\sum_{j=0}^{\infty}c_{j}{\hat{E}}^{j}\varphi(x,t).$ (30)
where $c_{j}$ are constants. Due to the degeneracy, it can be difficult to work with the general solution, but there is a specific selection of constants in the linear combination that can simplify the wave function. If the constants are written as
$c_{j}=\frac{1}{j!}\frac{\delta t^{j}}{(i\hbar)^{j}},$ (31)
the general solution can be written as
$\Psi(x,t)=\varphi(x,t-\delta t),$ (32)
where $\varphi(x,t-\delta t)=\psi(x,t)$ defined by Eq.(16). Even though one might assume that the general solution can be obtained straightforwardly by solving Eq.(12) and Eq.(14), it is important to note that this is just a particular case of the general solution. The possibility that different selections of constants $c_{j}$ will lead to different effects of the degeneracy must not be discarded.
Finally, an important property of this system can be deduced when the wave function in Eq.(32) is invariant under a unitary transformation of the operator Eq.(17), that is
${\hat{U}}_{x}\Psi(x,t)=\exp\left(i\frac{q{\cal E}}{\hbar}\delta x\delta t\right)\Psi(x,t),$ (33)
which leads to the condition that
$\frac{q{\cal E}}{\hbar}\delta x\delta t=2\pi n,\quad n\in\mathbb{N}.$ (34)
To interpret the above quantity, we rewrite it as follows
$\frac{q^{2}}{\hbar}{\cal E}\delta x\frac{\delta t}{q}=2\pi n,$ (35)
then we note that $q^{2}/\hbar$ is the inverse of Klitzing’s constant [16], ${\cal E}\delta x$ is the voltage, and $q/\delta t$ is the current of a single particle. Therefore, redefining Plank’s constant as $h=2\pi\hbar$ and rearranging
$R=\frac{h}{q^{2}}n,$ (36)
where $R$ is the resistance of the system. This means that the resistance produced by the particle is quantized as integer multiples of the von Klitzing constant.
IV Constant electromagnetic field
As mentioned in the introduction, another autonomous system with time-dependent conserved operators is the case of a constant electromagnetic field. The analysis for the situation where the electric field is perpendicular to the magnetic field can be found in reference [3]. Here, it will be consider the scenario where the electric field is parallel to the magnetic field. In this case, the same electric potential than in the previous section is used. On the other hand, to consider the magnetic field in Hamiltonian Eq.(1), it is necessary to define the vector potential. In this case ${\bf A}=B(0,0,y)$ satisfies the requirement that the magnetic field is parallel to the electric field. This selection of potentials leads to the following Hamiltonian
${\hat{H}}=\frac{1}{2m}\left\{{\hat{p}}_{x}^{2}+{\hat{p}}_{y}^{2}+({\hat{p}}_{z}-m\omega_{c}y)^{2}\right\}-q{\cal E}x$ (37)
where cyclotron frequency $\omega_{c}=qB/mc$ was defined. This hamiltonian has the following conserved operators
${\hat{\pi}_{x}}={\hat{p}}_{x}-q{\cal E}t,$ (38)
${\hat{\pi}_{y}}={\hat{p}}_{y}-m\omega_{c}z,$ (39)
${\hat{\pi}_{z}}={\hat{p}}_{z},$ (40)
and
${\hat{E}}=i\hbar\frac{\partial}{\partial t}.$ (41)
The Schrödinger equation can be written as
$i\hbar\frac{\partial\psi}{\partial t}={\hat{H}}\psi,$ (42)
where $\psi=\psi(x,y,z,t)$. For this specific situation, it is helpful to write down the wave function as the product of two time-dependent functions as follows
$\psi(x,y,z,t)=\phi_{1}(x,t)\phi_{2}(y,z,t),$ (43)
substituting in the Schrödinger Eq.(42), it leads to next couple of equations
$i\hbar\frac{\partial\phi_{1}}{\partial t}=\frac{1}{2m}{\hat{p}_{x}}^{2}\phi_{1}-q{\cal E}x\phi_{1},$ (44)
and
$i\hbar\frac{\partial\phi_{2}}{\partial t}=\frac{1}{2m}\left\{{\hat{p}}_{y}^{2}\phi_{2}+({\hat{p}}_{z}-m\omega_{c}y)^{2}\phi_{2}\right\}.$ (45)
The equation (44) is exactly the same than the one solved in the previous section, Eq.(9). Therefore, the previous analysis is valid and the wave function $\phi_{1}$ can be written as Eq.(23)
$\phi_{1}(x,t)=\frac{1}{L}\exp\left(-i\frac{q^{2}{\cal E}^{2}}{6m\hbar}t^{3}+i\frac{q{\cal E}}{\hbar}tx\right).$ (46)
On the other hand, the solutions for the Eq.(45) can be found with the help of the conserved operators Eq.(39) and Eq.(40). First, we proceed by separating the time coordinate as usual, by
$\phi_{2}(y,z,t)=\exp\left(-i\frac{E}{\hbar}t\right)\overline{\phi}_{2}(y,z),$ (47)
leading to the eigenvalue equation
$E\overline{\phi}_{2}=\frac{1}{2m}\left\{{\hat{p}}_{y}^{2}\overline{\phi}_{2}+({\hat{p}}_{z}-m\omega_{c}y)^{2}\overline{\phi}_{2}\right\},$ (48)
where $E$ is the energy of the system. Then, one can use these operators to write down a couple of eigenvalue equations
${\hat{p}}_{y}\overline{\phi}_{2}-m\omega_{c}z\overline{\phi}_{2}=m\omega_{c}\delta z\overline{\phi}_{2},$ (49)
${\hat{p}}_{z}\overline{\Phi}_{2}=m\omega_{c}\delta y\overline{\Phi}_{2},$ (50)
where $\delta z$ and $\delta y$ are constants. Both expressions will lead to valid solutions of the system. The simplest one is Eq.(50), it leads to a solution of the form
$\overline{\Phi}_{2}=\exp\left(i\frac{m\omega_{c}}{\hbar}z\delta y\right){\cal D}(y),$ (51)
where ${\cal D}(y)$ is a function to be determined by substituting in Eq.(48). Hence, this function satisfies the equation
$E{\cal D}=\frac{1}{2m}\left\{{\hat{p}}_{y}^{2}{\cal D}+m^{2}\omega_{c}^{2}(y-\delta y)^{2}{\cal D}\right\}.$ (52)
This equation can be identify as the displaced harmonic oscillator. Therefore,
${\cal D}(y)=\varphi_{n}\left(\sqrt{\frac{m\omega_{c}}{\hbar}}(y-\delta y)\right),$ (53)
where
$\varphi_{n}(\xi)=\frac{1}{\sqrt{2^{n}n!}}\left(\frac{m\omega_{c}}{\pi\hbar}\right)^{1/4}\exp(-\xi^{2})H_{n}(\xi),$ (54)
such that $H_{n}$ are the Hermite polynomials and the spectrum is given by the Landau’s levels
$E_{n}=\hbar\omega_{c}\left(n+\frac{1}{2}\right).$ (55)
This procedure is basically the application of Landau’s ansatz. However, there is an important mathematical detail to be considered. The expression Eq.(51), is a separable variable solution of the form $f(z)g(y)$, but if one attempts to formally separate the Eq.(48) it can be notice that in fact the equation can not be separated. This detail gives and insight that the solution of the Eq.(49) is still needed. From Eq.(49), one obtains that
$\overline{\phi}_{2}=\exp\left(i\frac{m\omega_{c}}{\hbar}y(z-\delta z)\right){\cal C}(z),$ (56)
where ${\cal C}(z)$ is a function to be determined . By substituting this last expression in Eq.(48) one obtains the following equation
$E{\cal C}=\frac{1}{2m}\left\{\hat{p}_{z}^{2}{\cal C}+m^{2}\omega_{c}^{2}(z-\delta z)^{2}{\cal C}\right\},$ (57)
once again, the displaced harmonic oscillator equation is obtained. Hence, the solutions are
${\cal C}(z)=\varphi_{n}\left(\sqrt{\frac{m\omega_{c}}{\hbar}}(z-\delta z)\right),$ (58)
where $\varphi_{n}$ is given by the expression Eq.(54), having the Landau levels as eigenvalues Eq.(55) .
At this point, is helpful to use the conserved operators to define a set of unitary operators
${\hat{U}}_{x}=\exp\left(-i\frac{\delta x}{\hbar}({\hat{p}}_{x}-q{\cal E}t)\right),$ (59)
${\hat{U}}_{y}=\exp\left(-i\frac{\delta y}{\hbar}({\hat{p}}_{y}-m\omega_{c}z)\right),$ (60)
${\hat{U}}_{z}=\exp\left(-i\frac{\delta y}{\hbar}{\hat{p}}_{z}\right),$ (61)
${\hat{U}}_{t}=\exp\left(i\frac{\delta t}{\hbar}{\hat{E}}\right).$ (62)
Such that they define the symmetries of the system
${\hat{H}}-{\hat{E}}={\hat{U}}_{i}^{\dagger}({\hat{H}}-{\hat{E}}){\hat{U}}_{i},\quad i=x,y,z,t.$ (63)
They can be used to rewrite the wave functions Eq.(51) and Eq.(56) as a unitary transformation of their respective harmonic oscillator, that is
$\overline{\Phi}_{2}={\hat{U}}_{y}\varphi_{n}\left(\sqrt{\frac{m\omega_{c}}{\hbar}}y\right),$ (64)
and
$\overline{\phi}_{2}={\hat{U}}_{z}\exp\left(i\frac{m\omega_{c}}{\hbar}yz\right)\varphi_{n}\left(\sqrt{\frac{m\omega_{c}}{\hbar}}z\right).$ (65)
This observation is useful to define a couple of time dependent wave functions using the respective inverse unitary transform
$\zeta_{n}(x,y,t)=\exp\left(-i\frac{E_{n}}{\hbar}t\right)\phi_{1}{\hat{U}}_{y}^{\dagger}\overline{\Phi}_{2},$ (66)
and
$\overline{\zeta}_{n}(x,y,z,t)=\exp\left(-i\frac{E_{n}}{\hbar}t\right)\phi_{1}{\hat{U}}_{z}^{\dagger}\overline{\phi}_{2},$ (67)
both of the above functions still have the Landau’s levels as eigenvalues of the Hamiltonian Eq.(37).
Before writing down the general solution, it is important to consider the degeneration of the system. At the beginning of this work the expression Eq.(5) was obtained for a time dependent operator. However, for the particular case when the operator is time-independent, for instance $\hat{\pi}_{y}$, the following can be written when the equality is applied to $\zeta_{n}$
${\hat{H}}(\hat{\pi}_{y}\zeta_{n})=\hat{\pi}_{y}({\hat{H}}\zeta_{n})=E_{n}(\hat{\pi}_{y}\zeta_{n}).$ (68)
Similarly, for the operator $\hat{\pi}_{z}$ and the function $\overline{\zeta}_{n}$ the following equality is obtained
${\hat{H}}(\hat{\pi}_{z}\overline{\zeta}_{n})=\hat{\pi}_{z}({\hat{H}}\overline{\zeta}_{n})=E_{n}(\hat{\pi}_{z}\overline{\zeta}_{n}).$ (69)
This last two equalities gives the insight that $\hat{\pi}_{y}\zeta_{n}$ and $\hat{\pi}_{z}\overline{\zeta}_{n}$ are also solutions of Hamiltonian Eq.(37) having the Landau’s levels as eigenvalues, even though if after the application of the respective operator the wave function obtained is not proportional to $\zeta_{n}$ or $\overline{\zeta}_{n}$. This observation can be easily generalized to any $j\in\mathbb{N}$ applications of the operators, which means that $\hat{\pi}_{y}^{j}\zeta_{n}$ and $\hat{\pi}_{z}^{j}\overline{\zeta}_{n}$ are also eigenfunctions of the system. On the other hand, the situation with the time-dependent operator $\hat{\pi}_{x}$ is not so different. Once again, from Eq.(5) applied to the function $\zeta_{n}$, which satisfies the time-dependent Schrödinger equation
$\hat{H}(\hat{\pi}_{x}\zeta_{n})=i\hbar\frac{\partial(\hat{\pi}_{x}\zeta_{n})}{\partial t},$ (70)
similarly for $\overline{\zeta}_{n}$
$\hat{H}(\hat{\pi}_{x}\overline{\zeta}_{n})=i\hbar\frac{\partial(\hat{\pi}_{x}\overline{\zeta}_{n})}{\partial t}.$ (71)
It can be noted that $\hat{\pi}_{x}\zeta_{n}$ and $\hat{\pi}_{x}\overline{\zeta}_{n}$ are also solutions of the Schrödinger equation. This observation can be generalized to $j$ applications of the same operators and one can realize that $\hat{\pi}_{x}^{j}\zeta_{n}$ and $\hat{\pi}_{x}^{j}\overline{\zeta}_{n}$ are also solutions of the Schrödinger equation. A final important observation is that the wave functions obtained due to this part of the degeneration are time dependent and since they are not proportional to the original functions $\zeta_{n}$ or $\overline{\zeta}_{n}$ the aplication of the operator $\hat{E}$ will also give a new solution of the Scrödinger equation and also will be the case for $j$ applications of the operator. Thus, the general solution, $\Psi$, is the linear combination of the solutions considering the degeneration of the system
$\Psi=\sum_{n,j,j^{\prime},k}a_{njj^{\prime}k}\hat{E}^{k}\hat{\pi}_{x}^{j}\hat{\pi}_{y}^{j^{\prime}}\zeta_{n}+\sum_{n,j,j^{\prime},k}\overline{a}_{njj^{\prime}k}\hat{E}^{k}\hat{\pi}_{x}^{j}\hat{\pi}_{z}^{j^{\prime}}\overline{\zeta}_{n},$ (72)
where $a_{njj^{\prime}k}$ and $\overline{a}_{njj^{\prime}k}$ are constants.
This last expression could be difficult to address analytically. However, there is a particular case for the constant which is worth to mention here. Being the constants
$a_{n,j,j^{\prime},k}=a_{n}\frac{1}{k!}\frac{\delta t^{k}}{(i\hbar)^{k}}\frac{(-1)}{j!}\frac{\delta x^{j}}{(i\hbar)^{j}}\frac{(-1)}{j^{\prime}!}\frac{\delta y^{j^{\prime}}}{(i\hbar)^{j^{\prime}}},$ (73)
and
$\overline{a}_{n,j,j^{\prime},k}=\overline{a}_{n}\frac{1}{k!}\frac{\delta t^{k}}{(i\hbar)^{k}}\frac{(-1)}{j!}\frac{\delta x^{j}}{(i\hbar)^{j}}\frac{(-1)}{j^{\prime}!}\frac{\delta y^{j^{\prime}}}{(i\hbar)^{j^{\prime}}},$ (74)
they lead to the following for of the general solution
$\Psi=\sum_{n}a_{n}{\hat{U}}_{t}{\hat{U}}_{x}{\hat{U}}_{y}\zeta_{n}+\sum_{n}\overline{a}_{n}{\hat{U}}_{t}{\hat{U}}_{x}{\hat{U}}_{z}\overline{\zeta}_{n}.$ (75)
If one propose this wave function to be invariant under the action of the unitary operator ${\hat{U}}_{x}$, meaning that ${\hat{U}}_{x}\Psi=\Psi$, one can realize that the following equality is obtained
${\hat{U}}_{x}\Psi=\exp\left(i\frac{q{\cal E}}{\hbar}\delta x\delta t\right)\Psi,$ (76)
implying, once again, that the invariance of the wave function under this unitary transformation implies the quantization of resistance Eq.(36).
V Conclusions
For both autonomous systems presented, it was shown that time-dependent conserved operators can be found and used to solve the Schrödinger equation. The degeneracy of the systems was analyzed, using the conserved operators as generators of solutions of the partial differential equation. Then, these operators were used to define the symmetries of the systems. It was proved that the invariance of the wave function under the action of the unitary operator, $\hat{U}_{x}$, leads to the quantization of resistance in integer multiples of Klitzing’s constant, in agreement with experimental results [4, 5, 6]. Finally, it is worth noting that Tao and Wu already pointed out that the degeneracy of the system is crucial for quantized resistance [20], a result that aligns with the findings presented in this work.
Acknowledgment
This work was supported by DGAPA-UNAM Posdoctoral Program (POSDOC).